# V-HOP: Visuo-Haptic 6D Object Pose Tracking

### v1

@@ 3.4 Training Paradigm
The rendered image is generated using $\widetilde{\mathbf{T}}$, while the object point hand is transformed based on $\widetilde{\mathbf{T}}$; in contrast, the RGB-D image and hand point cloud represent actual observations. The model estimates the relative pose $\Delta\widehat{\mathbf{T}}=(\Delta\widehat{\mathbf{R}},\Delta\widehat{\mathbf{t}})$ between the pose hypothesis and observation. The model is optimized using the $L_{2}$ loss:
@@
where we use quaternion representations for rotations. The estimated pose $\widehat{\mathbf{T}}=(\widehat{\mathbf{R}},\widehat{\mathbf{t}})$ is:
@@
We further incorporate an attractive loss ($\mathcal{L}_{a}$) and a penetration loss ($\mathcal{L}_{p}$) to encourage the object to make contact with the tactile point cloud $\mathcal{S}_{p,c}$ and avoid penetrating the hand point cloud $\mathcal{P}_{h}$. We first transform the initial hypothesized object pose cloud $\mathcal{P}_{o}$ using the estimated pose ${\widehat{\mathcal{P}}_{o}=\widehat{\mathbf{T}}\;\widetilde{\mathbf{T}}^{-1}\;\mathcal{P}_{o}}$, where $\mathcal{P}_{o}$ is in homogenous form.
@@
The penetration loss avoids penetrations between the object and the gripper Yang et al. (2021, 2024); Brahmbhatt et al. (2019):
@@
where $\mathbf{p}_{o}$ represents the nearest neighbor of each point $\mathbf{p}_{h}$ in the hand point cloud $\mathcal{P}_{h}$. Our overall loss is:
@@
where we set $\alpha=1$ and $\beta=0.001$ empirically. We optimize the model using the AdamW Loshchilov and Hutter (2018) optimizer with an initial learning rate of 0.0004 and train the model for 20 epochs.
4 Experiments
4.1 Multi-embodied Dataset
Existing visuo-haptic datasets were not publicly available Dikhale et al. (2022); Li et al. (2024b); Wan et al. (2024) at the time of completing this work and focused on a single gripper Suresh et al. (2024), leaving the question of generalization to novel embodiments unanswered. Consequently, we develop a multi-embodied dataset (Fig. 3) using NVIDIA Isaac Sim to enable cross-embodiment learning and thorough evaluation. Our dataset comprises approximately 1,550,000 images collected across eight grippers and thirteen objects. We utilize 85% of the data for training and the rest for validation. The camera trajectories are sampled on the semi-sphere around the gripper, which has a random radius between 0.5 and 2.5 meters. We selected graspable YCB object Calli et al. (2015) and grippers used in prior works Ding et al. (2024); Murrilo et al. (2024). Additional details about the dataset can be found in the appendix.
@@ 4.1 Multi-embodied Dataset
Table 1: Per-object comparison of AUC metrics for ADD and ADD-S. The row of novel object is grayed out. Both metrics are the higher, the better. The best results are bolded.
Hand Name AUC Metric ViTa FP V-HOP ADD 24.48 74.45 76.20 Allegro Hand ADD-S 77.60 88.74 90.48 ADD 24.63 77.67 79.06 Barrett Hand ADD-S 77.74 88.72 91.73 ADD 21.99 48.16 57.49 D’Claw (Unseen) ADD-S 76.00 77.06 85.48 ADD 24.56 70.22 70.15 Inspire Hand ADD-S 77.65 84.22 87.28 ADD 23.88 64.17 69.96 LEAP Hand ADD-S 77.55 83.06 88.05 ADD 23.40 79.48 79.14 Robotiq 3-Finger ADD-S 76.87 89.39 90.61 ADD 24.40 61.01 62.75 SCHUNK SVH ADD-S 76.74 78.58 82.96 ADD 23.81 58.77 60.27 Shadow Hand ADD-S 75.55 73.24 79.35 ADD $\uparrow$ 23.93 66.29 68.90 All ADD-S $\uparrow$ 76.87 82.37 86.62
Table 2: Per-gripper comparison of AUC metrics for ADD and ADD-S. Our dataset contains eight grippers. We train the model on seven grippers, leaving one gripper (D’Claw) unseen.
4.2 Pose Tracking Comparison
@@
  • Area under the curve (AUC) of ADD and ADD-S Hinterstoisser et al. (2013); Xiang et al. (2018), and
  • ADD(-S)-0.1d He et al. (2022): ADD/ADD-S that is less than 10% of the object diameter.
We compare V-HOP against the current state-of-the-art approaches in visual pose tracking (FoundationPose Wen et al. (2024), or FP in short) and visuo-tactile pose estimation (ViTa Dikhale et al. (2022)). To ensure a fair comparison, we finetune FoundationPose and train ViTa on our multi-embodied dataset. To verify the generalizability of the novel object and novel gripper, we exclude one object (pudding_box) and one gripper (D’Claw) during training.
@@
In Tab. 1, we show the performance for each object. V-HOP consistently outperforms ViTa and FoundationPose (FP) on most objects with respect to ADD and across all objects in terms of ADD-S. On average, our approach delivers an improvement of 4% in ADD and 5% in ADD-S compared to FoundationPose. Notably, V-HOP demonstrates strong performance on unseen objects, highlighting the potential of our model to generalize effectively to novel objects.
Similarly, Tab. 2 illustrates the performance of each gripper. In line with its object performance, V-HOP outperforms its counterparts on most grippers in terms of ADD and across all grippers in ADD-S. Moreover, V-HOP demonstrates robust performance on unseen grippers, further emphasizing the generalizability of our unified haptic representation.
@@
Table 3: Ablations of input modalities. Our results confirm the effectiveness of combining visual and haptic modalities.
@@
4.3 Ablation on Modalities
We conduct an ablation study on the input modalities to evaluate the effectiveness of the haptic representation. Specifically, we train two ablated versions of V-HOP: one without tactile feedback and another without visual input, as shown in Tab. 3. To exclude tactile input, we remove all “in contact” point labels (Equation 3). Our results indicate that visual input significantly contributes to performance, likely due to the richness of visual information, including texture and spatial details. This finding aligns with previous studies on human perception systems, which suggest that vision plays a dominant role in visuo-haptic integration Kassuba et al. (2013). Similarly, tactile feedback is crucial; without it, performance degrades notably because reasoning about hand-object contact during interactions becomes more difficult.
4.4 Occlusion’s Effect on the Performance
@@
4.5 Pose Tracking on FeelSight
Method GT Seg ADD-S $\downarrow$ ADD-S-0.1d $\uparrow$ FPS $\uparrow$ NeuralFeels Suresh et al. (2024) 2.14 98.95 3 V-HOP (Ours) 1.46 98.45 32
Table 4: Performance on the FeelSight Dataset. For consistency with the metric used in NeuralFeels Suresh et al. (2024), this experiment reports the direct ADD-S metric Xiang et al. (2018) (in mm) rather than the AUC of ADD-S used in other experiments.
To evaluate the generalizability of V-HOP, we benchmark it against NeuralFeels Suresh et al. (2024), a recently introduced optimization-based visuo-tactile pose tracking approach, using their proposed Feelsight dataset. Specifically, we focus on the occlusion subset of the dataset, FeelSight-Occlusion, which presents significant challenges due to severe occlusions. This subset requires robust generalization capabilities as it includes a novel embodiment (the Allegro hand equipped with DIGIT fingertips), a novel sensor type (a vision-based tactile sensor), and a novel object (a Rubik’s cube). For a fair comparison, we compare against their model-based tracking approach, which uses almost the same inputs as V-HOP but with the ground-truth segmentation mask (GT Seg).
The results are presented in Tab. 4. V-HOP achieves a 32% lower ADD-S error compared to NeuralFeels and has a similar ADD-S-0.1d score. It is important to note that NeuralFeels leverages the ground-truth segmentation mask, which helps in more accurate object localization, whereas V-HOP does not have such an input, further underscoring its robustness and adaptability.
@@ 4.5 Pose Tracking on FeelSight
5 Sim-to-Real Transfer Experiments
To validate the real-world effectiveness of our approach, we perform sim-to-real experiments using our robot platform (Fig. 1). Our bimanual platform comprises dual Franka Research 3 robotic arms Haddadin et al. (2022) and Barrett Hands BH8-282. Our Barrett Hand has 4 degrees of freedom (DoF) and 96 taxels: 24 taxels on each fingertip and 24 taxels on the palm. Each taxel comprises a capacitive cell capable of detecting forces within a range of 10 N/cm2 with a resolution of 0.01 N. For egocentric visual input, we use a MultiSense SLB RGB-D camera, which combines a MultiSense S7 stereo camera and a Hokuyo UTM-30LX-EW laser scanner. We utilize FoundationPose to provide the initial frame pose estimate and CNOS Nguyen et al. (2023); Kirillov et al. (2023) to provide the segmentation task.
@@ 5 Sim-to-Real Transfer Experiments
5.1 Pose Tracking Experiments
@@
5.2 Bimanual Handover Experiment
@@
    Use the left arm to grasp the object from the right hand and place it into a designated bin.
@@
Table 5: Success rate on bimanual handover task.
@@
V-HOP enables the motion planner to handle objects in random positions and adapt to dynamic scenarios, such as human perturbations. For instance, a human may move the object during task execution, remove it from the gripper, or reposition it on the table (Fig. 7). Due to the integration of haptic feedback, V-HOP accurately tracks the object’s pose, allowing the robot to promptly detect and respond to changes, such as the object leaving the gripper. On the contrary, FoundationPose loses tracking during handover or grasping failure (Fig. 6) and leads to collisions. In Tab. 5, we show the success rate for each object for five trials. V-HOP has 40% higher success rate on average compared to FoundationPose.
5.3 Can-in-Mug Experiment
@@
Table 6: Success rate on Can-in-Mug task.
@@
The Can-in-Mug task (Fig. 8) involves grasping a tomato can and inserting it into a mug. The bimanual version requires the robot to also grasp the mug and insert the can in the center. Successful execution hinges on precise pose estimation for both objects, as any noise in their poses can lead to failure. Our results (Tab. 6) demonstrate that V-HOP, by integrating visual and haptic inputs, delivers more stable tracking and a higher overall success rate.
5.4 Contribution of each modality
@@
Figure 9: Weights of visual and haptic modalities to the final prediction. We overlay the modality weights calculated using GradCAM Selvaraju et al. (2020) in the top-right corner.
@@
In this study, we examine the contribution of visual and haptic inputs to the final prediction. We adapt Grad-CAM Selvaraju et al. (2020), utilizing the final normalization layer of the Transformer encoder as the target layer. Figure 9 illustrates the weight distribution across the visual and haptic modalities. Our findings suggest that when the gripper is not in contact with an object, the model predominantly relies on visual inputs. However, as the gripper establishes contact and occlusion becomes more severe, the model increasingly shifts its reliance toward haptic inputs. This finding confirms the choice of self-attention mechanism to emulate human’s “optimal integration” principle.
6 Related Works
@@
In this work, we consider the problem of 6D object pose tracking problem, which has been widely studied as a visual problem Wen et al. (2020a); Li et al. (2018); Wen et al. (2024); Deng et al. (2021). In particular, we focus on model-based tracking approaches, which assume access to the object’s CAD model. While model-free approaches Wen and Bekris (2021); Wen et al. (2023); Suresh et al. (2024) exist, they fall outside the scope of this work. Visual pose tracking has achieved significant progress on established benchmarks, such as BOP Hodan et al. (2024). Despite these successes, deploying such systems in real-world robotic applications remains challenging, especially under scenarios with high occlusion and dynamic interactions, such as in-hand manipulation tasks.
@@
To address these challenges, prior research has explored combining visual and tactile information to improve pose tracking robustness Li et al. (2023); Suresh et al. (2024); Dikhale et al. (2022); Wan et al. (2024); Rezazadeh et al. (2023); Tu et al. (2023); Gao et al. (2023); Li et al. (2024b). These approaches leverage learning-based techniques to estimate object poses by fusing visuo-tactile inputs. However, these methods estimate poses on a per-frame basis, which lacks temporal coherence. Additionally, cross-embodiment and domain generalization remain significant hurdles, limiting their scalability and practicality for broad deployment.
@@
More recent works aim to overcome some of these limitations. For example, Liu et al. (2024) proposes an optimization-based approach that integrates tactile data with visual pose tracking using an ad-hoc slippage detector and velocity predictor. Suresh et al. (2024) extend the model-free tracking frameworks BundleTrack Wen and Bekris (2021) and BundleSDF Wen et al. (2023) by combining visual and tactile point clouds within a pose graph optimization framework. However, these approaches are only validated on a single embodiment and suffer from computational inefficiencies Suresh et al. (2024), which present challenges for real-time deployment in dynamic manipulation tasks.
7 Limitation

### v2

@@ III-D Training Paradigm
[…]
@@
[…]
@@
We further incorporate an attractive loss ($\mathcal{L}_{a}$) and a penetration loss ($\mathcal{L}_{p}$) to encourage the object to make contact with the tactile point cloud $\mathcal{S}_{p,c}$ and avoid penetrating the gripper point cloud $\mathcal{P}_{h}$. We first transform the initial hypothesized object pose cloud $\mathcal{P}_{o}$ using the estimated pose $\widehat{\mathcal{P}}_{o}=\widehat{\mathbf{T}}\;\widetilde{\mathbf{T}}^{-1}\;\mathcal{P}_{o}$}, where $\mathcal{P}_{o}$ is in homogenous form.
@@
The penetration loss avoids penetrations between the object and the gripper [76, 77, 4]:
@@
where $\mathbf{p}_{o}$ represents the nearest neighbor of each point $\mathbf{p}_{h}$ in the gripper point cloud $\mathcal{P}_{h}$. Our overall loss is:
@@
where we set $\alpha=1$ and $\beta=0.001$ empirically. We optimize the model using the AdamW [43] optimizer with an initial learning rate of 0.0004 and train the model for 20 epochs.
IV Experiments
IV-A Multi-embodied Dataset
Existing visuo-haptic datasets were not publicly available [8, 33, 61] at the time of completing this work and focused on a single gripper [54], leaving the question of generalization to novel embodiments unanswered. Consequently, we develop a multi-embodied dataset (Fig. 3) using NVIDIA Isaac Sim to enable cross-embodiment learning and thorough evaluation. Our dataset comprises approximately 1,550,000 images collected across eight grippers and thirteen objects. We utilize 85% of the data for training and the rest for validation. The camera trajectories are sampled on the semi-sphere around the gripper, which has a random radius between 0.5 and 2.5 meters. We selected graspable YCB object [5] and grippers used in prior works [9, 45]. Additional details about the dataset can be found in the appendix.
In this paper, we follow the sim-to-real paradigm and utilize only synthetic data for training. Increasing real-world training data could indeed help mitigate the sim-to-real gap. However, as demonstrated in recent work [70], leveraging a large-scale synthetic dataset enriched with domain randomization can yield superior real-world performance compared to small-scale real-world datasets. Our synthesized dataset exemplifies this principle and supports our robust real-world performance. Collecting real-world data with comparable scale and diversity would be both challenging and resource-intensive. Moreover, our unified haptic representation leverages point cloud representation to maintain invariance across various tactile sensors. Consequently, our sim-to-real experiments (Sec. V) demonstrate robust performance and eliminate the need for costly real-world data collection.
@@ IV-A Multi-embodied Dataset
TABLE I: Per-object comparison of AUC metrics for ADD and ADD-S. The row of novel object is grayed out. Both metrics are the higher, the better. The best results are bolded.
Gripper Name AUC Metric ViTa FP V-HOP ADD 24.48 74.45 76.20 Allegro Hand ADD-S 77.60 88.74 90.48 ADD 24.63 77.67 79.06 Barrett Hand ADD-S 77.74 88.72 91.73 ADD 21.99 48.16 57.49 D’Claw (Unseen) ADD-S 76.00 77.06 85.48 ADD 24.56 70.22 70.15 Inspire Hand ADD-S 77.65 84.22 87.28 ADD 23.88 64.17 69.96 LEAP Hand ADD-S 77.55 83.06 88.05 ADD 23.40 79.48 79.14 Robotiq 3-Finger ADD-S 76.87 89.39 90.61 ADD 24.40 61.01 62.75 SCHUNK SVH ADD-S 76.74 78.58 82.96 ADD 23.81 58.77 60.27 Shadow Hand ADD-S 75.55 73.24 79.35 ADD $\uparrow$ 23.93 66.29 68.90 All ADD-S $\uparrow$ 76.87 82.37 86.62
TABLE II: Per-gripper comparison of AUC metrics for ADD and ADD-S. Our dataset contains eight grippers. We train the model on seven grippers, leaving one gripper (D’Claw) unseen.
IV-B Pose Tracking Comparison
@@
  • Area under the curve (AUC) of ADD and ADD-S [21, 72], and
  • ADD(-S)-0.1d [18]: ADD/ADD-S that is less than 10% of the object diameter.
We compare V-HOP against the current state-of-the-art approaches in visual pose tracking (FoundationPose [70], or FP in short) and visuo-tactile pose estimation (ViTa [8]). To ensure a fair comparison, we finetune FoundationPose and train ViTa on our multi-embodied dataset. To verify the generalizability of the novel object and novel gripper, we exclude one object (pudding_box) and one gripper (D’Claw) during training.
@@
In Tab. I, we show the performance for each object. V-HOP consistently outperforms ViTa and FoundationPose (FP) on most objects with respect to ADD and across all objects in terms of ADD-S. On average, our approach delivers an improvement of 4% in ADD and 5% in ADD-S compared to FoundationPose. Notably, V-HOP demonstrates strong performance on unseen objects, highlighting the potential of our model to generalize effectively to novel objects.
Similarly, Tab. II illustrates the performance of each gripper. In line with its object performance, V-HOP outperforms its counterparts on most grippers in terms of ADD and across all grippers in ADD-S. Moreover, V-HOP demonstrates robust performance on unseen grippers, further emphasizing the generalizability of our unified haptic representation.
@@
TABLE III: Ablations of input modalities. Our results confirm the effectiveness of combining visual and haptic modalities.
Fusion Type AUC ADD ADD-0.1d AUC ADD-S ADD-S-0.1d Late Fusion 47.56 17.57 70.43 51.66 Early Fusion (Ours) 68.90 48.55 86.62 77.83
TABLE IV: Ablations of fusion strategies. We evaluate the performance of early fusion and late fusion strategies.
IV-C Ablation on Modalities
We conduct an ablation study on the input modalities to evaluate the effectiveness of the haptic representation. Specifically, we train two ablated versions of V-HOP: one without tactile feedback and another without visual input, as shown in Tab. III. To exclude tactile input, we remove all “in contact” point labels (Equation 3). Our results indicate that visual input significantly contributes to performance, likely due to the richness of visual information, including texture and spatial details. This finding aligns with previous studies on human perception systems, which suggest that vision plays a dominant role in visuo-haptic integration [24]. Similarly, tactile feedback is crucial; without it, performance degrades notably because reasoning about gripper-object contact during interactions becomes more difficult.
@@ IV-C Ablation on Modalities
IV-D Ablation on Fusion Strategies
We perform ablation studies on different modality fusion strategies: early fusion and late fusion. Early fusion refers to fusion at the input or feature level, the one we presented in Fig. 2. Late fusion strategy fuses the visual and tactile modalities at the result level, where each modality has a separate branch to estimate its result [59]. As shown in Tab. IV, the late fusion strategy results in an average ADD score of 47.56 and an ADD-S score of 70.43, which underperforms our early fusion design by 30.97% in ADD and 18.69% in ADD-S. The results confirm the necessity to fuse the visual and haptic modalities at the feature level.
@@ IV-D Ablation on Fusion Strategies
IV-E Occlusion’s Effect on the Performance
@@
IV-F Pose Tracking on FeelSight
Method GT Seg ADD-S $\downarrow$ ADD-S-0.1d $\uparrow$ FPS $\uparrow$ NeuralFeels [54] 2.14 98.95 3 V-HOP (Ours) 1.46 98.45 32
TABLE V: Performance on the FeelSight Dataset. For consistency with the metric used in NeuralFeels [54], this experiment reports the direct ADD-S metric [72] (in $m$) rather than the AUC of ADD-S used in other experiments.
To evaluate the generalizability of V-HOP, we benchmark it against NeuralFeels [54], a recently introduced optimization-based visuo-tactile pose tracking approach, using their proposed Feelsight dataset. Specifically, we focus on the occlusion subset of the dataset, FeelSight-Occlusion, which presents significant challenges due to severe occlusions. This subset requires robust generalization capabilities as it includes a novel embodiment (the Allegro hand equipped with DIGIT fingertips), a novel sensor type (a vision-based tactile sensor), and a novel object (a Rubik’s cube). For a fair comparison, we compare against their model-based tracking approach, which uses almost the same inputs as V-HOP but with the ground-truth segmentation mask (GT Seg).
The results are presented in Tab. V. V-HOP achieves a 32% lower ADD-S error compared to NeuralFeels and has a similar ADD-S-0.1d score. It is important to note that NeuralFeels leverages the ground-truth segmentation mask, which helps in more accurate object localization, whereas V-HOP does not have such an input, further underscoring its robustness and adaptability.
@@ IV-F Pose Tracking on FeelSight
V Sim-to-Real Transfer Experiments
To validate the real-world effectiveness of our approach, we perform sim-to-real experiments using our robot platform (Fig. 1). Our bimanual platform comprises dual Franka Research 3 robotic arms [17] and Barrett Hands BH8-282. Our Barrett Hand has 4 degrees of freedom (DoF) and 96 taxels: 24 taxels on each fingertip and 24 taxels on the palm. Each taxel comprises a capacitive cell capable of detecting forces within a range of 10 N/cm2 with a resolution of 0.01 N. For egocentric visual input, we use a MultiSense SLB RGB-D camera, which combines a MultiSense S7 stereo camera and a Hokuyo UTM-30LX-EW laser scanner. We utilize FoundationPose to provide the initial frame pose estimate and CNOS [47, 25] to provide the segmentation task.
V-A Pose Tracking Experiments
@@
V-B Bimanual Handover Experiment
@@
    Use the left arm to grasp the object from the right gripper and place it into a designated bin.
@@
TABLE VI: Success rate on bimanual handover task.
@@
V-HOP enables the motion planner to handle objects in random positions and adapt to dynamic scenarios, such as human perturbations. For instance, a human may move the object during task execution, remove it from the gripper, or reposition it on the table (Fig. 7). Due to the integration of haptic feedback, V-HOP accurately tracks the object’s pose, allowing the robot to promptly detect and respond to changes, such as the object leaving the gripper. On the contrary, FoundationPose loses tracking during handover or grasping failure (Fig. 6) and leads to collisions. In Tab. VI, we show the success rate for each object for five trials. V-HOP has 40% higher success rate on average compared to FoundationPose.
V-C Can-in-Mug Experiment
@@
TABLE VII: Success rate on Can-in-Mug task.
@@
The Can-in-Mug task (Fig. 8) involves grasping a tomato can and inserting it into a mug. The bimanual version requires the robot to also grasp the mug and insert the can in the center. Successful execution hinges on precise pose estimation for both objects, as any noise in their poses can lead to failure. Our results (Tab. VII) demonstrate that V-HOP, by integrating visual and haptic inputs, delivers more stable tracking and a higher overall success rate.
V-D Contribution of each modality
@@
Figure 9: Weights of visual and haptic modalities to the final prediction. We overlay the modality weights calculated using GradCAM [51] in the top-right corner.
@@
In this study, we examine the contribution of visual and haptic inputs to the final prediction. We adapt Grad-CAM [51], utilizing the final normalization layer of the Transformer encoder as the target layer. Figure 9 illustrates the weight distribution across the visual and haptic modalities. Our findings suggest that when the gripper is not in contact with an object, the model predominantly relies on visual inputs. However, as the gripper establishes contact and occlusion becomes more severe, the model increasingly shifts its reliance toward haptic inputs. This finding confirms the choice of self-attention mechanism to emulate human’s “optimal integration” principle.
VI Related Works
@@
In this work, we consider the problem of 6D object pose tracking problem, which has been widely studied as a visual problem [66, 35, 70, 7]. In particular, we focus on model-based tracking approaches, which assume access to the object’s CAD model. While model-free approaches [65, 69, 54] exist, they fall outside the scope of this work. Visual pose tracking has achieved significant progress on established benchmarks, such as BOP [22]. Despite these successes, deploying such systems in real-world robotic applications remains challenging, especially under scenarios with high occlusion and dynamic interactions, such as in-hand manipulation tasks.
@@
To address these challenges, prior research has explored combining visual and tactile information to improve pose tracking robustness [32, 54, 8, 61, 50, 59, 14, 33]. These approaches leverage learning-based techniques to estimate object poses by fusing visuo-tactile inputs. However, these methods estimate poses on a per-frame basis, which lacks temporal coherence. Additionally, cross-embodiment and domain generalization remain significant hurdles, limiting their scalability and practicality for broad deployment.
@@
More recent works aim to overcome some of these limitations. For example, Liu et al. [41] proposes an optimization-based approach that integrates tactile data with visual pose tracking using an ad-hoc slippage detector and velocity predictor. Suresh et al. [54] extend the model-free tracking frameworks BundleTrack [65] and BundleSDF [69] by combining visual and tactile point clouds within a pose graph optimization framework. However, these approaches are only validated on a single embodiment and suffer from computational inefficiencies [54], which present challenges for real-time deployment in dynamic manipulation tasks.
VII Limitation